# Robust Online Conformal Prediction under Uniform Label Noise

### v2

@@ B.4 Common non-conformity scores
Adaptive Prediction Set (APS). (Romano et al., 2020b) In the APS method, the non-conformity score of a data pair $(\bm{x},y)$ is calculated by accumulating the sorted softmax probability, defined as:
@@
Regularized Adaptive Prediction Set (RAPS). (Angelopoulos et al., 2021) The non-conformity score function of RAPS encourages a small set size by adding a penalty, as formally defined below:
@@
Sorted Adaptive Prediction Set (SAPS). (Huang et al., 2023) Recall that APS calculates the non-conformity score by accumulating the sorted softmax values in descending order. However, the softmax probabilities typically exhibit a long-tailed distribution, allowing for easy inclusion of those tail classes in the prediction sets. To alleviate this issue, SAPS discards all the probability values except for the maximum softmax probability when computing the non-conformity score. Formally, the non-conformity score of SAPS for a data pair $(\bm{x},y)$ can be calculated as
@@
B.5 Additional experiments for Section 4
To empirically validate Proposition 4.1, we compare the performance of standard online conformal prediction under different noise rates, with a ResNet18 model on CIFAR-100 and ImageNet datasets. We employ LAC score (Sadinle et al., 2019) to generate prediction sets, and use noisy labels to update the threshold with decaying learning rates $\eta_{t}=1/t^{1/2+\varepsilon}$ where $\varepsilon=0.1$. Figure 4 shows that label noise introduces a deviation between the achieved and target coverage rate $1-\alpha$, and larger noise rates would lead to more substantial coverage distortion.
(a) CIFAR-100
(b) ImageNet
Figure 4: Performance comparison of standard online conformal prediction under different noise rates, with a ResNet18 model on CIFAR-100 and ImageNet datasets.
B.6 Additional experiments on different non-conformity score functions
We evaluate NR-OCP (Ours) against the standard online conformal prediction (Baseline) that updates the threshold with noisy labels for both constant and dynamic learning rate schedules (see Eq. (2) and Eq. (5)). We use LAC (Table 3), APS (Table 4), RAPS (Table 5) and SAPS (Table 6) scores to generate prediction sets with error rates $\alpha\in\{0.1,0.05\}$, and employ noise rates $\epsilon\in\{0.05,0.1,0.15\}$. A detailed description of these non-conformity scores is provided in Appendix B.4.
@@ B.6 Additional experiments on different non-conformity score functions
Constant $\epsilon=0.05$ Baseline 2.744 1.900 31.61 56.84 2.747 1.601 364.5 599.1
@@
Dynamic $\epsilon=0.05$ Baseline 3.978 2.833 11.54 41.75 4.333 2.889 87.27 391.8
@@
Constant $\epsilon=0.05$ Baseline 4.556 2.933 5.862 21.17 4.728 3.508 13.35 79.88
@@
Dynamic $\epsilon=0.05$ Baseline 4.000 2.433 4.693 13.03 4.157 2.122 11.31 31.38
@@
Constant $\epsilon=0.05$ Baseline 3.978 3.533 5.275 26.71 4.716 4.509 13.73 164.7
@@
Dynamic $\epsilon=0.05$ Baseline 4.322 3.254 4.883 19.42 4.642 3.457 12.57 60.94
@@
Constant $\epsilon=0.05$ Baseline 4.489 3.100 4.982 17.92 4.462 3.537 12.53 83.04
@@
Dynamic $\epsilon=0.05$ Baseline 4.122 2.600 4.992 13.68 4.859 1.866 13.46 28.28
@@
B.7 Additional experiments on different model architectures
We evaluate NR-OCP (Ours) against the standard online conformal prediction (Baseline) that updates the threshold with noisy labels for both constant and dynamic learning rate schedules (see Eq. (2) and Eq. (5)), employing ResNet50 (Table 7), DenseNet121 (Table 8), and VGG16 (Table 9). We use LAC scores to generate prediction sets with error rates $\alpha\in\{0.1,0.05\}$, and employ noise rates $\epsilon\in\{0.05,0.1,0.15\}$.
@@ B.7 Additional experiments on different model architectures
Constant $\epsilon=0.05$ Baseline 3.244 1.783 28.58 56.08 3.226 1.982 294.4 575.7
@@
Dynamic $\epsilon=0.05$ Baseline 4.166 2.724 10.91 41.95 4.397 3.104 44.86 361.8
@@
Constant $\epsilon=0.05$ Baseline 3.978 3.533 5.275 26.71 4.716 4.509 13.73 164.7
@@
Dynamic $\epsilon=0.05$ Baseline 4.322 3.254 4.883 19.42 4.642 3.457 12.57 60.94
@@
Constant $\epsilon=0.05$ Baseline 1.911 1.033 51.13 73.03 3.164 1.995 297.8 565.3
@@
Dynamic $\epsilon=0.05$ Baseline 2.678 1.388 40.96 66.81 4.608 3.044 51.64 374.1
@@
Appendix C Proof for Section 3
C.1 Proof for Proposition 3.1
Lemma C.1.
@@ Appendix C Proof for Section 3
Proposition C.2 (Restatement of Proposition 3.1).
@@
$\alpha-\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}\leq\frac{2-\epsilon}{1-\epsilon}\sqrt{2\log\left(\frac{4}{\delta}\right)}\cdot\frac{1}{\sqrt{T}}+\frac{1}{1-\epsilon}\left(\frac{1}{\eta}+1-\alpha\right)\cdot\frac{1}{T}+\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\right).$
@@
Part (a): Lemma G.3 gives us that
@@
$\mathbb{E}_{S}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},S_{t})\right]$ =$=\mathbb{P}\left\{S_{t}\leq\hat{\tau}_{t}\right\}-(1-\alpha)$
=$\overset{(a)}{=}\left(\frac{1}{1-\epsilon}\mathbb{P}\left\{\tilde{S}_{t}\leq\hat{\tau}_{t}\right\}-\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}\mathbb{P}\left\{S_{t,y}\leq\hat{\tau}_{t}\right\}\right)-(1-\alpha)$
@@
where (a) comes from Lemma G.1. In addition, Lemma G.3 implies that
@@
$\leq\frac{1}{1-\epsilon}\left|\sum_{t=1}^{T}\mathbb{E}_{\tilde{S}}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},\tilde{S})\right]-\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},\tilde{S})\right|+\frac{1}{1-\epsilon}\left|\sum_{t=1}^{T}\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},\tilde{S})\right|+\frac{\epsilon}{1-\epsilon}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\right)$
@@
where (a) follows from the assumption that $\hat{\tau}_{1}\in[0,1]$ and Lemma C.1. Thus, we have
@@
$\quad\sum_{t=1}^{T}\nabla_{\tau_{t}}l_{1-\alpha}\left(\tau_{t},S_{t}\right)$ $\leq\sqrt{2T\log\left(\frac{4}{\delta}\right)}+\frac{1}{1-\epsilon}\sqrt{2T\log\left(\frac{4}{\delta}\right)}+\frac{1}{1-\epsilon}\left(\frac{1}{\eta}+1-\alpha\right)+\frac{\epsilon}{1-\epsilon}\sum_{t=1}^{T}\left(\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]-(1-\alpha)\right)$
@@
$\alpha-\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}\leq\frac{2-\epsilon}{1-\epsilon}\sqrt{2\log\left(\frac{4}{\delta}\right)}\cdot\frac{1}{\sqrt{T}}+\frac{1}{1-\epsilon}\left(\frac{1}{\eta}+1-\alpha\right)\cdot\frac{1}{T}+\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\right)$
C.2 Proof for Proposition 3.2
Proposition C.3 (Restatement of Proposition 3.2).
@@
[…]
@@
[…]
@@
[…]
@@
$\quad\mathbb{E}_{\tilde{S},S_{y}}\bigg{[}(1-\alpha)\tau\left(\frac{1}{1-\epsilon}\mathds{1}\left\{\tilde{S}>\tau\right\}+\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}\mathds{1}\left\{S_{y}>\tau\right\}\right)+(1-\alpha)\left(\frac{\tilde{S}}{1-\epsilon}\mathds{1}\left\{\tilde{S}>\tau\right\}+\frac{\epsilon S_{y}}{K(1-\epsilon)}\sum_{y=1}^{K}\mathds{1}\left\{S_{y}>\tau\right\}\right)\bigg{]}$
@@
C.3 Proof for Proposition 3.3
Lemma C.4.
@@
$-\alpha\eta-\frac{\epsilon\eta}{1-\epsilon}\leq\hat{\tau}_{T}\leq 1-\alpha\eta+\frac{\eta}{1-\epsilon}.$ (8)
@@
where (a) follows from Lemma G.2.
@@
Proposition C.5 (Restatement of Proposition 3.3).
@@
where (a) follows from the assumption that $\hat{\tau}_{1}\in[0,1]$ and Lemma C.4. In addition, Lemma G.5 gives us that
@@
C.4 Proof for Proposition 3.4
Proposition C.6 (Restatement of Proposition 3.4).
@@
Lemma G.3 gives us that
@@
holds with at least $21-\delta/$ 2 probability. In addition, recall that

### v3

@@ B.4 Common non-conformity scores
Adaptive Prediction Set (APS). [34] In the APS method, the non-conformity score of a data pair $(\bm{x},y)$ is calculated by accumulating the sorted softmax probability, defined as:
@@
Regularized Adaptive Prediction Set (RAPS). [35] The non-conformity score function of RAPS encourages a small set size by adding a penalty, as formally defined below:
@@
Sorted Adaptive Prediction Set (SAPS). [36] Recall that APS calculates the non-conformity score by accumulating the sorted softmax values in descending order. However, the softmax probabilities typically exhibit a long-tailed distribution, allowing for easy inclusion of those tail classes in the prediction sets. To alleviate this issue, SAPS discards all the probability values except for the maximum softmax probability when computing the non-conformity score. Formally, the non-conformity score of SAPS for a data pair $(\bm{x},y)$ can be calculated as
@@
B.5 Additional experiments on different non-conformity score functions
We evaluate NR-OCP (Ours) against the standard online conformal prediction (Baseline) that updates the threshold with noisy labels for both constant and dynamic learning rate schedules (see Eq. (3)). We use LAC (Table 2), APS (Table 3), RAPS (Table 4) and SAPS (Table 5) scores to generate prediction sets with error rates $\alpha\in\{0.1,0.05\}$, and employ noise rates $\epsilon\in\{0.05,0.1,0.15\}$. A detailed description of these non-conformity scores is provided in Appendix B.4.
@@ B.5 Additional experiments on different non-conformity score functions
Constant
$\epsilon=0.05$ Baseline 2.744 1.900 31.61 56.84 2.747 1.601 364.5 599.1
@@
Dynamic
$\epsilon=0.05$ Baseline 3.978 2.833 11.54 41.75 4.333 2.889 87.27 391.8
@@
Constant
$\epsilon=0.05$ Baseline 4.556 2.933 5.862 21.17 4.728 3.508 13.35 79.88
@@
Dynamic
$\epsilon=0.05$ Baseline 4.000 2.433 4.693 13.03 4.157 2.122 11.31 31.38
@@
Constant
$\epsilon=0.05$ Baseline 3.978 3.533 5.275 26.71 4.716 4.509 13.73 164.7
@@
Dynamic
$\epsilon=0.05$ Baseline 4.322 3.254 4.883 19.42 4.642 3.457 12.57 60.94
@@
Constant
$\epsilon=0.05$ Baseline 4.489 3.100 4.982 17.92 4.462 3.537 12.53 83.04
@@
Dynamic
$\epsilon=0.05$ Baseline 4.122 2.600 4.992 13.68 4.859 1.866 13.46 28.28
@@
B.6 Additional experiments on different model architectures
We evaluate NR-OCP (Ours) against the standard online conformal prediction (Baseline) that updates the threshold with noisy labels for both constant and dynamic learning rate schedules (see Eq. (3) and Eq. (LABEL:eq:changing_noise_update_rule)), employing ResNet50 (Table 6), DenseNet121 (Table 7), and VGG16 (Table 8). We use LAC scores to generate prediction sets with error rates $\alpha\in\{0.1,0.05\}$, and employ noise rates $\epsilon\in\{0.05,0.1,0.15\}$.
@@ B.6 Additional experiments on different model architectures
Constant
$\epsilon=0.05$ Baseline 3.244 1.783 28.58 56.08 3.226 1.982 294.4 575.7
@@
Dynamic
$\epsilon=0.05$ Baseline 4.166 2.724 10.91 41.95 4.397 3.104 44.86 361.8
@@
Constant
$\epsilon=0.05$ Baseline 3.978 3.533 5.275 26.71 4.716 4.509 13.73 164.7
@@
Dynamic
$\epsilon=0.05$ Baseline 4.322 3.254 4.883 19.42 4.642 3.457 12.57 60.94
@@
Constant
$\epsilon=0.05$ Baseline 1.911 1.033 51.13 73.03 3.164 1.995 297.8 565.3
@@
Dynamic
$\epsilon=0.05$ Baseline 2.678 1.388 40.96 66.81 4.608 3.044 51.64 374.1
@@
Appendix C The impact of misestimated noise rate on NR-OCP
In this section, we conduct an additional experiment where the noise rate is unknown and needs to be estimated. In particular, we apply an existing algorithm [61] to estimate the noise rate without access to clean data during the training of the pre-trained model. Then, we employ this estimated noise rate in our method for online conformal prediction. The experiments are conducted on CIFAR-10 dataset, using ResNet18 model with noise rates $\epsilon\in\{0.1,0.2,0.3\}$. We employ LAC score to generate prediction sets and use a constant learning rate $\eta=0.05$. The results demonstrate that our method consistently achieves much lower coverage gaps and smaller prediction sets compared to the baseline. This highlights the robustness and applicability of our method even when the noise rate is not precisely estimated.
Table 9:
Noise rate Estimated Noise rate Method CovGap (%) Size
$\epsilon=0.1$ $\hat{\epsilon}=0.09$ Baseline 6.99 2.47
Ours 0.82 0.94
$\epsilon=0.2$ $\hat{\epsilon}=0.23$ Baseline 8.72 5.41
Ours 1.67 0.98
$\epsilon=0.3$ $\hat{\epsilon}=0.26$ Baseline 9.04 6.72
Ours 2.32 1.15
Appendix D Additional information on Strongly adaptive online conformal prediction
D.1 Strongly adaptive online conformal prediction
The SAOCP algorithm leverages techniques for minimizing the strongly adaptive regret [67] to perform online conformal prediction. In particular, SAOCP is a meta-algorithm that manages multiple experts, where each expert is itself an arbitrary online learning algorithm taking charge of its own active interval that has a finite lifetime. At each $t\in[T]$, SAOCP instantiates a new expert $\mathcal{A}_{t}$ with active interval [t,t+L(t)1][t,t+L(t)-1], where $L(t)$ is the lifetime:
$L(t)=g\cdot\max_{n\in\mathbb{Z}_{\geq 1}}\{2^{n}:t\equiv 0\pmod{2^{n}}\}$, (7)
where $g\in\mathbb{Z}_{\geq 1}$ is a multiplier for the lifetime of each expert. It can be shown that at most $g\log_{2}t$ experts are active at any time $t$ under the choice of $L(t)$ in (7), resulting in a total runtime of $O(T\log T)$ for SAOCP when $g=\Theta(1)$. At each time $t$, the threshold $\hat{\tau}_{t}$ is obtained by aggregating the predictions of the active experts.
$\hat{\tau}_{t}=\sum_{i\in\text{Active}(t)}p_{i,t}\hat{\tau}_{t}^{i}$
where the weight $\{p_{i,t}\}_{i\in[t]}$ is computed by the coin betting scheme [67, 68].
Choice of expert.
SAOCP employs Scale-Free OGD (dubbed SF-OGD) [68] as its expert. In particular, SF-OGD is a variant of OGD that decays its effective learning rate based on cumulative past gradient norms.
D.2 Noise-robust strongly adaptive online conformal prediction
In the following, we present a pseudo-algorithm of NR-SAOCP. The essence of this method is to update the threshold with our robust pinball loss. In Algorithm 1, we demonstrate how NR-SAOCP manages multiple experts to update the threshold. In Algorithm 2, we present how each expert updates the corresponding threshold.
Algorithm 1 Noise-Robust Strongly Adaptive Online Conformal Prediction (NR-SAOCP)
0: Target coverage $1-\alpha\in(0,1)$, initial threshold $\hat{\tau}_{0}$.
1:for $t=1,\ldots,T$ do
2:  Initialize new expert $\mathcal{A}_{t}=\text{NR-SF-OGD}(\alpha\leftarrow\alpha;\eta\leftarrow 1/\sqrt{3};\hat{\tau}_{1}\leftarrow\hat{\tau}_{t-1})$ (Algorithm 2), and set weight $w_{t,t}=0$
3:  Compute active set $\text{Active}(t)=\{i\in[T]:t-L(i)\leq i\leq t\},$ where $L(i)$ is defined in Eq. (7)
[…]
5:  Update the threshold $\hat{\tau}_{t}=\sum_{i\in\text{Active}(t)}p_{i}\hat{\tau}_{t}^{i}$
6:  Observe input $X_{t}$ and construct prediction set $\hat{C}_{t}(X_{t})$ as in Eq. (1)
7:  Observe true label $Y_{t}\in\mathcal{Y}$, compute non-conformity score $S_{t}=\mathcal{S}(X_{t},Y_{t})$ and $S_{t,y}=\mathcal{S}(X_{t},y)$ for all $y\in\mathcal{Y}$
8:  for $i\in\text{Active}(t)$ do
9:   Update expert $A_{i}$ with robust pinball loss and obtain next predicted radius $\hat{\tau}_{t+1}^{i}$
10:   Compute
$g_{i,t}=\begin{cases}\tilde{l}(\hat{\tau}_{t},S_{t},\{S_{t,y}\}_{y=1}^{K})-\tilde{l}(\hat{\tau}_{t}^{i},S_{t},\{S_{t,y}\}_{y=1}^{K})&\text{if }w_{t,i}>0\\ [\tilde{l}(\hat{\tau}_{t},S_{t},\{S_{t,y}\}_{y=1}^{K})-\tilde{l}(\hat{\tau}_{t}^{i},S_{t},\{S_{t,y}\}_{y=1}^{K})]_{+}&\text{if }w_{t,i}\leq 0\end{cases}$
where $\tilde{l}$ is the robust pinball loss defined in Eq. (4).
11:   Update expert weight $w_{i,t+1}=\frac{1}{t-i+1}\left(\sum_{j=i}^{t}g_{i,j}\right)\left(1+\sum_{j=i}^{t}w_{i,j}g_{i,j}\right)$
12:  end for
13:end for
Algorithm 2 Noise-Robust Scale-Free Online Gradient Descent (NR-SF-OGD)
[…]
1:for $t\geq 1$ do
2:  Observe input $X_{t}$ and construct prediction set $\hat{C}_{t}(X_{t})$ as in Eq. (1)
3:  Observe true label $Y_{t}\in\mathcal{Y}$, compute non-conformity score $S_{t}=\mathcal{S}(X_{t},Y_{t})$ and $S_{t,y}=\mathcal{S}(X_{t},y)$ for all $y\in\mathcal{Y}$
4:  Update the threshold:
$\hat{\tau}_{t+1}=\hat{\tau}_{t}-\eta\,\frac{\nabla\tilde{l}_{\hat{\tau}_{t}}(\hat{\tau}_{t},S_{t},\{S_{t,y}\}_{y=1}^{K})}{\sqrt{\sum_{i=1}^{t}\|\nabla_{\hat{\tau}_{i}}\tilde{l}(\hat{\tau}_{i},S_{i},\{S_{i,y}\}_{y=1}^{K})\|_{2}^{2}}}$
where $\tilde{l}$ is the robust pinball loss defined in Eq. (4).
5:end for
Appendix E Omitted proofs
E.1 Proof for Proposition 3.1
Lemma E.1.
@@ Appendix E Omitted proofs
Proposition E.2 (Restatement of Proposition 3.1).
@@
$\alpha-\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}\leq$ $\frac{2-\epsilon}{1-\epsilon}\sqrt{2\log\left(\frac{4}{\delta}\right)}\cdot\frac{1}{\sqrt{T}}+\frac{1}{1-\epsilon}\left(\frac{1}{\eta}+1-\alpha\right)\cdot\frac{1}{T}$
$+\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\right).$
@@
Part (a): Lemma F.3 gives us that
@@
$\mathbb{E}_{S}\left[\nabla_{\hat{\tau}_{t}}l_{1-\alpha}(\hat{\tau}_{t},S_{t})\right]$ $=\mathbb{P}\left\{S_{t}\leq\hat{\tau}_{t}\right\}-(1-\alpha)$
$\overset{(a)}{=}\left(\frac{1}{1-\epsilon}\mathbb{P}\left\{\tilde{S}_{t}\leq\hat{\tau}_{t}\right\}-\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}\mathbb{P}\left\{S_{t,y}\leq\hat{\tau}_{t}\right\}\right)-(1-\alpha)$
@@
where (a) comes from Lemma F.1. In addition, Lemma F.3 implies that
@@
[…]
$\quad+\frac{\epsilon}{1-\epsilon}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\right)$
@@
where (a) follows from the assumption that $\hat{\tau}_{1}\in[0,1]$ and Lemma E.1. Thus, we have
@@
$\quad\sum_{t=1}^{T}\nabla_{\tau_{t}}l_{1-\alpha}\left(\tau_{t},S_{t}\right)$ $\leq\sqrt{2T\log\left(\frac{4}{\delta}\right)}+\frac{1}{1-\epsilon}\sqrt{2T\log\left(\frac{4}{\delta}\right)}+\frac{1}{1-\epsilon}\left(\frac{1}{\eta}+1-\alpha\right)$
$\quad+\frac{\epsilon}{1-\epsilon}\sum_{t=1}^{T}\left(\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]-(1-\alpha)\right)$
@@
$\alpha-\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\left\{Y_{t}\notin\mathcal{C}_{t}(X_{t})\right\}$ $\leq\frac{2-\epsilon}{1-\epsilon}\sqrt{2\log\left(\frac{4}{\delta}\right)}\cdot\frac{1}{\sqrt{T}}+\frac{1}{1-\epsilon}\left(\frac{1}{\eta}+1-\alpha\right)\cdot\frac{1}{T}$
$\quad+\frac{\epsilon}{1-\epsilon}\cdot\frac{1}{T}\sum_{t=1}^{T}\left((1-\alpha)-\frac{1}{K}\mathbb{E}\left[\mathcal{C}_{t}(X_{t})\right]\right)$
E.2 Proof for Proposition 4.1
Proposition E.3 (Restatement of Proposition 4.1).
@@
$=\alpha\tau\mathbb{P}\left\{\tau\geq S\right\}-\alpha\int_{0}^{\tau}s\mathbb{P}\left\{S=s\right\}ds-(1-\alpha)\tau\mathbb{P}\left\{\tau<S\right\}+(1-\alpha)\int_{\tau}^{1}s\mathbb{P}\left\{S=s\right\}ds$
$=\underbrace{\alpha\tau\left(\frac{1}{1-\epsilon}\mathbb{P}\left\{\tilde{S}\leq\tau\right\}+\frac{\epsilon}{1-\epsilon}\mathbb{P}\left\{\bar{S}\leq\tau\right\}\right)}_{(a)}-\underbrace{\alpha\int_{0}^{\tau}s\left(\frac{1}{1-\epsilon}\mathbb{P}\left\{\tilde{S}\leq\tau\right\}+\frac{\epsilon}{1-\epsilon}\mathbb{P}\left\{\bar{S}\leq\tau\right\}\right)ds}_{(b)}$-
$\quad\underbrace{(1-\alpha)\tau\left(\frac{1}{1-\epsilon}\mathbb{P}\left\{\tilde{S}>\tau\right\}+\frac{\epsilon}{1-\epsilon}\mathbb{P}\left\{\bar{S}>\tau\right\}\right)}_{(c)}+\underbrace{(1-\alpha)\int_{\tau}^{1}s\left(\frac{1}{1-\epsilon}\mathbb{P}\left\{\tilde{S}>\tau\right\}+\frac{\epsilon}{1-\epsilon}\mathbb{P}\left\{\bar{S}>\tau\right\}\right)ds}_{(d)}$
@@
$\alpha\int_{0}^{\tau}s\left(\frac{1}{1-\epsilon}\mathbb{P}\left\{\tilde{S}\leq\tau\right\}+\frac{\epsilon}{1-\epsilon}\mathbb{P}\left\{\bar{S}\leq\tau\right\}\right)ds$ $=\alpha\int_{0}^{\tau}s\left(\frac{1}{1-\epsilon}\mathbb{P}\left\{\tilde{S}\leq\tau\right\}+\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}\mathbb{P}\left\{S_{y}\leq\tau\right\}\right)ds$
@@
$\quad(1-\alpha)\int_{\tau}^{1}s\left(\frac{1}{1-\epsilon}\mathbb{P}\left\{\tilde{S}>\tau\right\}+\frac{\epsilon}{1-\epsilon}\mathbb{P}\left\{\bar{S}>\tau\right\}\right)ds$
$=(1-\alpha)\int_{0}^{\tau}s\left(\frac{1}{1-\epsilon}\mathbb{P}\left\{\tilde{S}>\tau\right\}+\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}\mathbb{P}\left\{S_{y}>\tau\right\}\right)ds$
@@
$\quad\mathbb{E}_{\tilde{S},S_{y}}\bigg[(1-\alpha)\tau\left(\frac{1}{1-\epsilon}\mathds{1}\left\{\tilde{S}>\tau\right\}+\frac{\epsilon}{K(1-\epsilon)}\sum_{y=1}^{K}\mathds{1}\left\{S_{y}>\tau\right\}\right)$
$\quad+(1-\alpha)\left(\frac{\tilde{S}}{1-\epsilon}\mathds{1}\left\{\tilde{S}>\tau\right\}+\frac{\epsilon S_{y}}{K(1-\epsilon)}\sum_{y=1}^{K}\mathds{1}\left\{S_{y}>\tau\right\}\right)\bigg]$
@@
E.3 Proof for Proposition 4.2
Lemma E.4.
@@
$\alpha\eta-\frac{\epsilon\eta}{1-\epsilon}\leq\hat{\tau}_{T}\leq 1-\alpha\eta+\frac{\eta}{1-\epsilon}.$ (9)
@@
where (a) follows from Lemma F.2.
@@
Proposition E.5 (Restatement of Proposition 4.2).
@@
where (a) follows from the assumption that $\hat{\tau}_{1}\in[0,1]$ and Lemma E.4. In addition, Lemma F.5 gives us that
@@
E.4 Proof for Proposition 4.3
Proposition E.6 (Restatement of Proposition 4.3).
@@
Lemma F.3 gives us that
@@
holds with at least $1-\delta/2$ probability. In addition, recall that